# Consistency of Neural Causal Partial Identification

### v1

@@ Appendix C Proof of Consistency
C.1 Inconsistent Counterexample (Proposition 2)
Proposition 5.
There exists a constant $c>0$ and an SCM $\mathcal{M}^{*}$ satisfying Assumptions 1-5 with a backdoor causal graph such that there is no confounding in $\mathcal{M}^{*}$. For any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ with the same causal graph such that $W(P^{\mathcal{M}^{*}}(\boldsymbol{V}),P^{\mathcal{M}_{\epsilon}}(\boldsymbol{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
@@ C.1 Inconsistent Counterexample (Proposition 2)
Let $\boldsymbol{V}=\{X,T,Y\}$ be the covariate, treatment and outcome respectively. $T$ is a binary variable. Let structure equations of causal model $\mathcal{M}_{\delta}$ be
@@
It is easy to see that $\mathcal{M}^{*}$ satisfies Assumption 1-5. Some calculation gives $W(P^{\mathcal{M}^{*}}(\boldsymbol{V}),P^{\mathcal{M}_{\delta}}(\boldsymbol{V}))\leqslant\delta/2$. It is easy to calculate the ATE of the two models.
@@
This example implies that even as the Wasserstein distance between $P^{\mathcal{M}^{*}}(\boldsymbol{V}),P^{\mathcal{M}_{\delta}}(\boldsymbol{V})$ converges to zero, their ATEs do not change. As mentioned in the main body, this problem is caused by the violation of Lipschitz continuity assumption for $\mathcal{M}_{\delta}$. Note that in $\mathcal{M}_{\delta}$, $\mathbb{E}[Y|X,T]=X/\delta$. As $\delta\rightarrow 0$, the Lipschitz constant explodes.
This problem may also arise in (6). If the distribution ball $B_{n}=\{\hat{\mathcal{M}}:W(P^{\hat{\mathcal{M}}},P^{\mathcal{M}^{*}}_{n})\leqslant\alpha_{n}\}$ includes a small ball around the true distribution $S_{\epsilon}=\{\hat{\mathcal{M}}:W(P^{\hat{\mathcal{M}}},P^{\mathcal{M}^{*}})\leqslant\epsilon\}\subset B_{n}$ and the NCM is expressive enough to approximate all the SCMs in the $S_{\epsilon}$, this example tells us the confidence interval may not shrink to one point as sample size increases to infinity even if the model is identifiable.
C.2 Proof of Theorem 4
To prove Theorem 4, we will need the following proposition.
Proposition 6.
@@ C.2 Proof of Theorem 4
    $\|g_{n}(\theta_{1})-g_{n}(\theta_{2})\|\leqslant L_{g}d(\theta_{1},\theta_{2})+\tau_{n},\quad\forall\theta_{1},\theta_{2}\in\Theta$ (21)
@@
    There exists a compact subset $\tilde{\Theta}_{\infty}\subset\Theta_{\infty}$ such that for any $\theta_{0}\in\tilde{\Theta}_{\infty}$, there exists $\theta\in\Theta_{n}$ such that $d(\theta,\theta_{0})\leqslant\epsilon_{n}.$
@@
$s.t.$ $\ g_{n}(\theta)\leqslant L_{g}\epsilon_{n}+\delta_{n},$ (22)
@@
$\min_{\theta\in\tilde{\Theta}_{\infty}}$ $f(\theta),$
$s.t.$ $\ g(\theta)=0.$ (24)
We assume that the feasible region of (23) and (24) are nonempty. Let $f_{n}^{*},f^{*},\tilde{f}^{*}$ be the minimal of (22), (23) and (24) respectively. Then, $[\liminf_{k\rightarrow\infty}f_{n}^{*},\limsup_{k\rightarrow\infty}f_{n}^{*}]\subset[f^{*},\tilde{f}^{*}]$.
Proof of Proposition 6.
By compactness of $\Theta_{\infty}$ and $\tilde{\Theta}_{\infty}$, the minimizers of (23) and (24) are achievable insider these two sets. Let $\theta^{*}\in\Theta_{\infty},\tilde{\theta}^{*}\in\tilde{\Theta}_{\infty}$ be the minimizer of (23) and (24). We first prove that $\limsup_{n\rightarrow\infty}f_{n}^{*}\leqslant\tilde{f}^{*}$. Note that by (21), the limiting function $g$ is $L_{g}$-Lipschitz. By condition 3, there exist $\theta_{n}\in\Theta_{n}$ such that $d\left(\theta_{n},\tilde{\theta}^{*}\right)\leqslant\epsilon_{n}$. Note that
@@
Therefore, $\theta_{n}$ is a feasible point of (22). We have $\limsup_{n\rightarrow\infty}f_{n}^{*}\leqslant\limsup_{n\rightarrow\infty}f({\theta}_{n})=f\left(\tilde{\theta}^{*}\right)=\tilde{f}^{*}$.
Next, we argue that $\liminf_{n\rightarrow\infty}f_{n}^{*}\geqslant f^{*}$. If this equation does not hold, there exists $\epsilon>0$ and subsequence $\left\{f_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ such that $f_{n_{k}}^{*}<f^{*}-\epsilon,\forall k\in\mathbb{N}$. By compactness of $\Theta_{\infty}$, for each $k$, there exists a subsequence of $\theta_{n_{k}}^{*}\in\Theta_{\infty}$ such that $\theta_{n_{k}}^{*}$ satisfies constraint of (22) and $f_{n_{k}}^{*}=f\left(\theta_{n_{k}}^{*}\right)$. By compactness, $\left\{\theta_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ has a converging subseqence. Without loss of generality, we may assume that $\left\{\theta_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ converges to $\hat{\theta}^{*}\in\Theta_{\infty}$. Since $g_{k}$ converge uniformly to $g$,
$\lim_{k\rightarrow\infty}g_{k}\left(\theta_{n_{k}}^{*}\right)\leqslant\lim_{k\rightarrow\infty}|g_{k}\left(\theta_{n_{k}}^{*}\right)-g\left(\theta_{n_{k}}^{*}\right)|+g\left(\theta_{n_{k}}^{*}\right)=g\left(\hat{\theta}^{*}\right)\leqslant\lim_{k\rightarrow\infty}\alpha_{k}=0.$
$\hat{\theta}^{*}$ is a feasible point of Equation 23. Since $f$ is continuous on $\Theta$,
$f\left(\hat{\theta}^{*}\right)=\lim_{k\rightarrow\infty}f\left(\theta_{n_{k}}^{*}\right)\leqslant f^{*}-\epsilon.$
@@
Proof of Theorem 4.
We begin by defining a proper metric space. By 3, all random variables are bounded. Suppose that $\max_{i,j}\{\|V_{i}\|_{\infty},\|U_{j}\|_{\infty}\}\leqslant K$. A canonical causal model $\mathcal{M}\in\mathcal{M}(\mathcal{G},\mathcal{F},\boldsymbol{U})$ is decided by
@@
where $f_{i}$ are functions in the structure equations (3) and $U_{j}$ are uniform parts of the latent variables. We denote $\mathcal{M}^{\theta}$ to be the SCM represented by $\theta$, the underlying SCM be $\mathcal{M}^{\theta^{*}}$ and $P^{\theta}$ to be the distribution of $\mathcal{M}^{\theta}$. We consider the space
@@
Theorem 1 states that the Wasserstein distance between two causal models is Lipschitz with respect to metric $d$. Now, we define $\Theta_{n}$. Let
$\mathcal{P}_{n}=\left\{{P}({\boldsymbol{U}}):\boldsymbol{U}\ \text{is the latent distribution of}\ \hat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})\right\}.$
In other words, $\mathcal{P}_{n}$ contains all the push-forward measures of the uniform distribution by neural networks. We denote $\Theta_{n}=\hat{\mathcal{F}}_{V_{1},n}\times\cdots\times\hat{\mathcal{F}}_{V_{n_{V}},n}\times\times\mathcal{P}_{n}$, where
^$\hat{\mathcal{F}}_{V_{i},n}=\begin{cases}\mathcal{F}_{0,n},&V_{i}\ \text{continuous},\\ \{f/\|f\|_{1}:f\in\mathcal{F}_{0,n}\}&V_{i}\ \text{categorical},\end{cases}$
and $\mathcal{F}_{0,n}$ is defined in Theorem 4. Note that by construction, latent variables are independent and $\mathcal{P}_{n}$ can be decomposed into direct produce $\mathcal{P}_{n,1}\times\cdots\times\mathcal{P}_{n,n_{U}}$. Let
$\Theta_{\infty}=\overline{\cup_{n=1}^{\infty}\Theta_{n}},g_{n}(\theta)=S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right),f(\theta)=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{\theta}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$
and $\tilde{\Theta}_{\infty}=\{\theta^{*}\}$. Note that $f(\theta)$ is a continuous function since by Theorem 1 and the fact that $F$ is Lispchitz continuous,
$|f(\theta)-f(\theta^{\prime})|$ $\leqslant\mathbb{E}_{t\sim\mu_{T}}|W({P}^{\theta}(\boldsymbol{V}(t)),{P}^{\theta^{\prime}}(\boldsymbol{V}(t))|$
@@
Now, we verify the conditions in Proposition 6.
@@
    $|g_{n}(\theta)-g_{n}(\theta^{\prime})|$ $=|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta^{\prime}}(\boldsymbol{V})\right)|$
    $\leqslant|W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta^{\prime}}(\boldsymbol{V})\right)|+4(\log(mn)+1)\lambda_{n}$
    $\leqslant W\left(P_{m_{n}}^{\theta}(\boldsymbol{V}),P_{m_{n}}^{\theta^{\prime}}(\boldsymbol{V})\right)+4(\log(mn)+1)\lambda_{n}$
@@
    where we use Lemma 7 in the second inequality and triangle inequality and Theorem 1 in the third inequality. By the condition in Theorem 4, the second term $(1+\log(nm_{n}))\lambda_{n}\rightarrow 0$ as $n\rightarrow\infty$. Therefore, (21) is verified with $\tau_{n}=4(1+\log(nm_{n}))\lambda_{n}$.
@@
    $|g_{n}(\theta)-g(\theta)|$ $=|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)|$
    $\leqslant|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)|$
    $\ \ \ \ \ \ \ \ +|W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)|.$
    By Lemma 7, $|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)|\leqslant 2(\log(mn)+1)\lambda_{n}$ and we get
    $|g_{n}(\theta)-g(\theta)|$ $\leqslant 2(\log(mn)+1)\lambda_{n}+|W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P_{m_{n}}^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)|$
    $\quad\ \ \ \ +|W\left(P^{\theta^{*}}_{n}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)-W\left(P^{\theta^{*}}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)|$
    $\leqslant 2(\log(mn)+1)\lambda_{n}+W\left(P_{m_{n}}^{\theta}(\boldsymbol{V}),P^{\theta}(\boldsymbol{V})\right)+W(P^{\theta^{*}}_{n}(\boldsymbol{V}),P^{\theta^{*}}(\boldsymbol{V}))$
    $\leqslant 2(\log(mn)+1)\lambda_{n}+O\left(W(P^{\theta^{*}}_{n}(\boldsymbol{U}),P^{\theta^{*}}(\boldsymbol{U}))+W\left(P^{\theta}(\boldsymbol{U}),P^{\theta}_{m_{n}}(\boldsymbol{U})\right)\right),$ (25)
    where we use Theorem 1 in the last inequality.
    We first bound $\sup_{\theta}W\left(P^{\theta}(\boldsymbol{U}),P_{m_{n}}^{\theta}(\boldsymbol{U})\right)$ using standard VC dimension argument. Note that $\{U_{1},\cdots,U_{n_{U}}\}$ are independent. By Lemma 5,
    $W\left(P^{\theta}(\boldsymbol{U}),P_{m_{n}}^{\theta}(\boldsymbol{U})\right)\leqslant\sum_{i=1}^{n_{U}}W\left(P^{\theta}(U_{i}),P_{m_{n}}^{\theta}(U_{i})\right).$ (26)
    By the construction in Theorem 3, $P^{\theta}(U_{i})\sim\sum_{j\in[N_{C,i}]}p_{j}P(f_{\theta_{i,j}}(Z_{i,j})),Z_{i,j}\sim\text{Unif}(0,1)$, where $f_{\theta_{i,j}}$ are neural networks with constant width and depth $L_{1,n}+L_{2,n}$, $N_{C,i}$ is the number of connected components of $\text{supp}(P(U_{i}))$ and $\theta_{i,j}$ are the parameters of the neural networks. By Lemma 4,
@@
    By [6], the pseudo-dimension of $\mathcal{NN}(\Theta(1),L)$ is $O\left(L^{2}\log L\right)$. By boundness of all the neural networks and Lemma 8  [^1] Note that we truncate the neural network to ensure boundness in Theorem 4. This extra truncate operation $\tilde{f}=\max\{-K,\min\{K,f\}\}$ can be viewed as two extra ReLu layers, so Lemma 8 is still applicable. , with probability at least
@@
    $\sup_{\theta}W\left(P^{\theta}(U_{i}),P_{m_{n}}^{\theta}(U_{i})\right)$ $\leqslant O\left(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}\right),$ (27)
@@
    Therefore, $\sum_{n=1}^{\infty}P(W(P^{\theta^{*}}({U}_{i}),P^{\theta^{*}}_{n}({U}_{i}))>\delta_{n})<\infty.$ By Borel-Cantelli lemma, with probability 1, there exist $N>0$ such that $W(P^{\theta^{*}}({U}_{i}),P^{\theta^{*}}_{n}({U}_{i}))\leqslant\delta_{n},\forall n>N$. By Equation 26, with probability 1, there exist $N>0$ such that
    $W(P^{\theta^{*}}(\boldsymbol{U}),P^{\theta^{*}}_{n}(\boldsymbol{U}))\leqslant O(\delta_{n}),\forall n>N.$ (28)
    Therefore, by Equations 25, 27 and 28, almost surely, the following inequality holds eventually,
@@
    where $s_{n}$ is defined in Theorem 4.

### v2

@@ Appendix C Proof of Consistency
C.1 Inconsistent Counterexample (Proposition 4.1)
Proposition C.1.
There exists a constant $c>0$ and an SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.6 with a backdoor causal graph such that there is no unobserved confounding in $\mathcal{M}^{*}$. For any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ with the same causal graph such that $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
@@
Let $\bm{V}=\{X,T,Y\}$ be the covariate, treatment and outcome respectively. $T$ is a binary variable. Let structure equations of causal model $\mathcal{M}_{\delta}$ be
@@
It is easy to see that $\mathcal{M}^{*}$ satisfies Assumption 2.3-3.6. Some calculation gives $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\delta}}(\bm{V}))\leqslant\delta/2$. It is easy to calculate the ATE of the two models.
@@
This example implies that even as the Wasserstein distance between $P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\delta}}(\bm{V})$ converges to zero, their ATEs do not change. As mentioned in the main body, this problem is caused by the violation of Lipschitz continuity assumption for $\mathcal{M}_{\delta}$. Note that in $\mathcal{M}_{\delta}$, $\mathbb{E}[Y|X,T]=X/\delta$. As $\delta\rightarrow 0$, the Lipschitz constant explodes.
This problem may also arise in (4.1). If the distribution ball $B_{n}=\{\widehat{\mathcal{M}}:W(P^{\widehat{\mathcal{M}}},P^{\mathcal{M}^{*}}_{n})\leqslant\alpha_{n}\}$ includes a small ball around the true distribution $S_{\epsilon}=\{\widehat{\mathcal{M}}:W(P^{\widehat{\mathcal{M}}},P^{\mathcal{M}^{*}})\leqslant\epsilon\}\subset B_{n}$ and the NCM is expressive enough to approximate all the SCMs in the $S_{\epsilon}$, this example tells us the confidence interval may not shrink to one point as sample size increases to infinity even if the model is identifiable.
C.2 Proof of Theorem 4.2
To prove Theorem 4.2, we will need the following proposition.
Proposition C.2.
@@
    $\|g_{n}(\theta_{1})-g_{n}(\theta_{2})\|\leqslant L_{g}d(\theta_{1},\theta_{2})+\tau_{n},\quad\forall\theta_{1},\theta_{2}\in\Theta$ (C.1)
@@
    There exists a compact subset $\widetilde{\Theta}_{\infty}\subset\Theta_{\infty}$ such that for any $\theta_{0}\in\widetilde{\Theta}_{\infty}$, there exists $\theta\in\Theta_{n}$ such that $d(\theta,\theta_{0})\leqslant\epsilon_{n}.$
@@
$s.t.$ $\ g_{n}(\theta)\leqslant L_{g}\epsilon_{n}+\delta_{n},$ (C.2)
@@
$\min_{\theta\in\widetilde{\Theta}_{\infty}}$ $f(\theta),$
$s.t.$ $\ g(\theta)=0.$ (C.4)
We assume that the feasible region of (C.3) and (C.4) are nonempty. Let $f_{n}^{*},f^{*},\widetilde{f}^{*}$ be the minimal of (C.2), (C.3) and (C.4) respectively. Then, $[\liminf_{k\rightarrow\infty}f_{n}^{*},\limsup_{k\rightarrow\infty}f_{n}^{*}]\subset[f^{*},\widetilde{f}^{*}]$.
Proof of Proposition C.2.
By compactness of $\Theta_{\infty}$ and $\widetilde{\Theta}_{\infty}$, the minimizers of (C.3) and (C.4) are achievable insider these two sets. Let $\theta^{*}\in\Theta_{\infty},\widetilde{\theta}^{*}\in\widetilde{\Theta}_{\infty}$ be the minimizer of (C.3) and (C.4). We first prove that $\limsup_{n\rightarrow\infty}f_{n}^{*}\leqslant\widetilde{f}^{*}$. Note that by (C.1), the limiting function $g$ is $L_{g}$-Lipschitz. By condition 3, there exist $\theta_{n}\in\Theta_{n}$ such that $d\left(\theta_{n},\widetilde{\theta}^{*}\right)\leqslant\epsilon_{n}$. Note that
@@
Therefore, $\theta_{n}$ is a feasible point of (C.2). We have $\limsup_{n\rightarrow\infty}f_{n}^{*}\leqslant\limsup_{n\rightarrow\infty}f({\theta}_{n})=f\left(\widetilde{\theta}^{*}\right)=\widetilde{f}^{*}$.
Next, we argue that $\liminf_{n\rightarrow\infty}f_{n}^{*}\geqslant f^{*}$. If this equation does not hold, there exists $\epsilon>0$ and subsequence $\left\{f_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ such that $f_{n_{k}}^{*}<f^{*}-\epsilon,\forall k\in\mathbb{N}$. By compactness of $\Theta_{\infty}$, for each $k$, there exists a subsequence of $\theta_{n_{k}}^{*}\in\Theta_{\infty}$ such that $\theta_{n_{k}}^{*}$ satisfies constraint of (C.2) and $f_{n_{k}}^{*}=f\left(\theta_{n_{k}}^{*}\right)$. By compactness, $\left\{\theta_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ has a converging subseqence. Without loss of generality, we may assume that $\left\{\theta_{n_{k}}^{*}\right\}_{k\in\mathbb{N}}$ converges to $\widehat{\theta}^{*}\in\Theta_{\infty}$. Since $g_{k}$ converge uniformly to $g$,
$\lim_{k\rightarrow\infty}g_{k}\left(\theta_{n_{k}}^{*}\right)\leqslant\lim_{k\rightarrow\infty}|g_{k}\left(\theta_{n_{k}}^{*}\right)-g\left(\theta_{n_{k}}^{*}\right)|+g\left(\theta_{n_{k}}^{*}\right)=g\left(\widehat{\theta}^{*}\right)\leqslant\lim_{k\rightarrow\infty}\alpha_{k}=0.$
$\widehat{\theta}^{*}$ is a feasible point of Equation C.3. Since $f$ is continuous on $\Theta$,
$f\left(\widehat{\theta}^{*}\right)=\lim_{k\rightarrow\infty}f\left(\theta_{n_{k}}^{*}\right)\leqslant f^{*}-\epsilon.$
@@
Proof of Theorem 4.2.
We begin by defining a proper metric space. By 3.2, all random variables are bounded. Suppose that $\max_{i,j}\{\|V_{i}\|_{\infty},\|U_{j}\|_{\infty}\}\leqslant K$. A canonical causal model $\mathcal{M}\in\mathcal{M}(\mathcal{G},\mathcal{F},\bm{U})$ is decided by
@@
where $f_{i}$ are functions in the structure equations (3.1) and $U_{j}$ are uniform parts of the latent variables. We denote $\mathcal{M}^{\theta}$ to be the SCM represented by $\theta$, the underlying SCM be $\mathcal{M}^{\theta^{*}}$ and $P^{\theta}$ to be the distribution of $\mathcal{M}^{\theta}$. We consider the space
@@
Theorem 3.3 states that the Wasserstein distance between two causal models is Lipschitz with respect to metric $d$. Now, we define $\Theta_{n}$. Let
$\mathcal{P}_{n}=\left\{{P}({\bm{U}}):\bm{U}\ \text{is the latent distribution of}\ \widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})\right\}.$
In other words, $\mathcal{P}_{n}$ contains all the push-forward measures of the uniform distribution by neural networks. We denote $\Theta_{n}=\widehat{\mathcal{F}}_{V_{1},n}\times\cdots\times\widehat{\mathcal{F}}_{V_{n_{V}},n}\times\times\mathcal{P}_{n}$, where
^$\widehat{\mathcal{F}}_{V_{i},n}=\begin{cases}\mathcal{F}_{0,n},&V_{i}\ \text{continuous},\\ \{f/\|f\|_{1}:f\in\mathcal{F}_{0,n}\}&V_{i}\ \text{categorical},\end{cases}$
and $\mathcal{F}_{0,n}$ is defined in Theorem 4.2. Note that by construction, latent variables are independent and $\mathcal{P}_{n}$ can be decomposed into direct produce $\mathcal{P}_{n,1}\times\cdots\times\mathcal{P}_{n,n_{U}}$. Let
$\Theta_{\infty}=\overline{\cup_{n=1}^{\infty}\Theta_{n}},g_{n}(\theta)=S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right),f(\theta)=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{\theta}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$
and $\widetilde{\Theta}_{\infty}=\{\theta^{*}\}$. Note that $f(\theta)$ is a continuous function since by Theorem 3.3 and the fact that $F$ is Lispchitz continuous,
$|f(\theta)-f(\theta^{\prime})|$ $\leqslant\mathbb{E}_{t\sim\mu_{T}}|W({P}^{\theta}(\bm{V}(t)),{P}^{\theta^{\prime}}(\bm{V}(t))|$
@@
Now, we verify the conditions in Proposition C.2.
@@
    $|g_{n}(\theta)-g_{n}(\theta^{\prime})|$ $=|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta^{\prime}}(\bm{V})\right)|$
    $\leqslant|W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta^{\prime}}(\bm{V})\right)|+4(\log(mn)+1)\lambda_{n}$
    $\leqslant W\left(P_{m_{n}}^{\theta}(\bm{V}),P_{m_{n}}^{\theta^{\prime}}(\bm{V})\right)+4(\log(mn)+1)\lambda_{n}$
@@
    where we use Lemma E.5 in the second inequality and triangle inequality and Theorem 3.3 in the third inequality. By the condition in Theorem 4.2, the second term $(1+\log(nm_{n}))\lambda_{n}\rightarrow 0$ as $n\rightarrow\infty$. Therefore, (C.1) is verified with $\tau_{n}=4(1+\log(nm_{n}))\lambda_{n}$.
@@
    $|g_{n}(\theta)-g(\theta)|$ $=|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}(\bm{V}),P^{\theta}(\bm{V})\right)|$
    $\leqslant|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)|$
    $\ \ \ \ \ \ \ \ +|W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}(\bm{V}),P^{\theta}(\bm{V})\right)|.$
    By Lemma E.5, $|S_{\lambda_{n}}\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)|\leqslant 2(\log(mn)+1)\lambda_{n}$ and we get
    $|g_{n}(\theta)-g(\theta)|$ $\leqslant 2(\log(mn)+1)\lambda_{n}+|W\left(P^{\theta^{*}}_{n}(\bm{V}),P_{m_{n}}^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}_{n}(\bm{V}),P^{\theta}(\bm{V})\right)|$
    $\quad\ \ \ \ +|W\left(P^{\theta^{*}}_{n}(\bm{V}),P^{\theta}(\bm{V})\right)-W\left(P^{\theta^{*}}(\bm{V}),P^{\theta}(\bm{V})\right)|$
    $\leqslant 2(\log(mn)+1)\lambda_{n}+W\left(P_{m_{n}}^{\theta}(\bm{V}),P^{\theta}(\bm{V})\right)+W(P^{\theta^{*}}_{n}(\bm{V}),P^{\theta^{*}}(\bm{V}))$
    $\leqslant 2(\log(mn)+1)\lambda_{n}+O\left(W(P^{\theta^{*}}_{n}(\bm{U}),P^{\theta^{*}}(\bm{U}))+W\left(P^{\theta}(\bm{U}),P^{\theta}_{m_{n}}(\bm{U})\right)\right),$ (C.5)
    where we use Theorem 3.3 in the last inequality.
    We first bound $\sup_{\theta}W\left(P^{\theta}(\bm{U}),P_{m_{n}}^{\theta}(\bm{U})\right)$ using standard VC dimension argument. Note that $\{U_{1},\cdots,U_{n_{U}}\}$ are independent. By Lemma E.3,
    $W\left(P^{\theta}(\bm{U}),P_{m_{n}}^{\theta}(\bm{U})\right)\leqslant\sum_{i=1}^{n_{U}}W\left(P^{\theta}(U_{i}),P_{m_{n}}^{\theta}(U_{i})\right).$ (C.6)
    By the construction in Theorem 3.8, $P^{\theta}(U_{i})\sim\sum_{j\in[N_{C,i}]}p_{j}P(f_{\theta_{i,j}}(Z_{i,j})),Z_{i,j}\sim\text{Unif}(0,1)$, where $f_{\theta_{i,j}}$ are neural networks with constant width and depth $L_{1,n}+L_{2,n}$, $N_{C,i}$ is the number of connected components of $\text{supp}(P(U_{i}))$ and $\theta_{i,j}$ are the parameters of the neural networks. By Lemma E.2,
@@
    By [6], the pseudo-dimension of $\mathcal{NN}(\Theta(1),L)$ is $O\left(L^{2}\log L\right)$. By boundness of all the neural networks and Lemma E.6  [^5] Note that we truncate the neural network to ensure boundness in Theorem 4.2. This extra truncate operation $\widetilde{f}=\max\{-K,\min\{K,f\}\}$ can be viewed as two extra ReLu layers, so Lemma E.6 is still applicable. , with probability at least
@@
    $\sup_{\theta}W\left(P^{\theta}(U_{i}),P_{m_{n}}^{\theta}(U_{i})\right)$ $\leqslant O\left(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}\right),$ (C.7)
@@
    Therefore, $\sum_{n=1}^{\infty}P(W(P^{\theta^{*}}({U}_{i}),P^{\theta^{*}}_{n}({U}_{i}))>\delta_{n})<\infty.$ By Borel-Cantelli lemma, with probability 1, there exist $N>0$ such that $W(P^{\theta^{*}}({U}_{i}),P^{\theta^{*}}_{n}({U}_{i}))\leqslant\delta_{n},\forall n>N$. By Equation C.6, with probability 1, there exist $N>0$ such that
    $W(P^{\theta^{*}}(\bm{U}),P^{\theta^{*}}_{n}(\bm{U}))\leqslant O(\delta_{n}),\forall n>N.$ (C.8)
    Therefore, by Equations C.5, C.7 and C.8, almost surely, the following inequality holds eventually,
@@
    where $s_{n}$ is defined in Theorem 4.2.